OpenStudy (anonymous):

please help!! log(2x+1)=1-logx

OpenStudy (anonymous):

I did part of it and got stumped...

OpenStudy (anonymous):

log (2x+1) + log x = 1 log x(2x+1) = 1 x(2x+1) = 10

OpenStudy (anonymous):

log(2x+1)=1-logx log(2x+1)=log10-logx log(2x+1)=log(10/x) 2x+1=10/x 2x^2+x=10 2x^2+x-10=0

OpenStudy (anonymous):

Same thing as divu is getting...

OpenStudy (anonymous):

2x^2+x-10=0 (2x+5)( x-2)=0

OpenStudy (anonymous):

Go , you got it from here !

OpenStudy (anonymous):

Thank you Basil!

OpenStudy (anonymous):

It's my pleasure :)

OpenStudy (anonymous):

Any my first day on this website;)

OpenStudy (anonymous):

And...

OpenStudy (anonymous):

Math wiz dude! :)

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

I'm just 14 though...